OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = –5 when x = 15, find y when x = –1.

OpenStudy (anonymous):

@hba @Orwell84 @Orwell84 @phi @eliassaab @ajprincess @XxNina26xX @bookworm00981 @phi

OpenStudy (anonymous):

@Kilorion @Orwell84 @bookworm00981 @ganeshie8 @Mashy @skullpatrol @ajprincess @Orwell84 @phi

OpenStudy (anonymous):

Example: x = -40 direct variation y = kx to find k, use y=2 when x = - 4 2 = k (-4), so k = 2/(-4) = -1/2 now y = (-1/2) x and when y = 20, 20 = (-1/2) x, so x = (-2)20 = -40